OpenStudy (anonymous):

Solve: 2(x + 1) = −4

OpenStudy (anonymous):

Distribute the 2 to the (x+1)

OpenStudy (anonymous):

x + 1 = -2 x = -3

OpenStudy (anonymous):

Dont Distribute Divide both sides by 2

OpenStudy (anonymous):

It Will be more easy....

OpenStudy (anonymous):

divide by 2 on both sides 2(x+1)/2 = -4/2 x+1=-2 subtract by 1 x+1-1=-2-1 x=-3